# Scalable Sobolev IPM for Probability Measures on a Graph

### v1

@@ Abstract
We investigate the Sobolev IPM problem for probability measures supported on a graph metric space. Sobolev IPM is an important instance of integral probability metrics (IPM), and is obtained by constraining a critic function within a unit ball defined by the Sobolev norm. In particular, it has been used to compare probability measures and is crucial for several theoretical works in machine learning. However, to our knowledge, there are no efficient algorithmic approaches to compute Sobolev IPM effectively, which hinders its practical applications. In this work, we establish a relation between Sobolev norm and weighted $L^{p}$-norm, and leverage it to propose a novel regularization for Sobolev IPM. By exploiting the graph structure, we demonstrate that the regularized Sobolev IPM provides a closed-form expression for fast computation. This advancement addresses long-standing computational challenges, and paves the way to apply Sobolev IPM for practical applications, even in large-scale settings. Additionally, the regularized Sobolev IPM is negative definite. Utilizing this property, we design positive-definite kernels upon the regularized Sobolev IPM, and provide preliminary evidences of their advantages on document classification and topological data analysis for measures on a graph.
@@ 1 Introduction
Probability measures: equal contribution are widely used to represent objects of interest across various research fields. For instance, in natural language processing, documents can be viewed as distributions over words (Sparck Jones, 1972; Kusner et al., 2015), or distributions over topics (Blei et al., 2003; Yurochkin et al., 2019). In computer vision and graphics, 3D objects are often represented as point clouds, which are distributions of 3D data points (Achlioptas et al., 2018; Hua et al., 2018; Wang et al., 2019; Wu et al., 2019). In topological data analysis (TDA), persistence diagrams (PD) are used to represent complex structural objects (Rieck et al., 2019; Zhao & Wang, 2019; Divol & Lacombe, 2021). Specifically, PDs can be regarded as distributions of 2D data points, where each point characterizes the birth and death time of a particular topological feature (Edelsbrunner & Harer, 2008).
@@
Our contributions are three-fold:
@@
  • We demonstrate that the regularized Sobolev IPM is negative definite. We then design positive definite kernels based on the regularized Sobolev IPM and provide preliminary evidence of their advantages in document classification and topological data analysis (TDA).
@@
In §2, we introduce the notations used throughout our proposals. Section §3 details our novel regularization for Sobolev IPM. In §4, we prove the metric property for the regularized Sobolev IPM and establish its connection to the original Sobolev IPM, ST, and OT on a graph. Additionally, we demonstrate the negative definiteness for the regularized Sobolev IPM and propose positive definite kernels based on it. Section §5 discusses related works. In §6, we empirically illustrate the computational advantages of the regularized Sobolev IPM and provide preliminary evidence of its benefits in document classification and TDA. Finally, §7 offers concluding remarks. Proofs for theoretical results and additional materials are deferred to the Appendices.
@@ 2 Preliminaries
Graph. We follow the same graph setting in Le et al. (2022) for measures and functions. Precisely, let $V,E$ denote the sets of nodes and edges respectively. We consider a connected, undirected, and physical [^1] In the sense that $V$ is a subset of Euclidean space $\mathbb{R}^{n}$; each edge $e\in E$ is the standard line segment in $\mathbb{R}^{n}$ connecting the two vertices of the edge $e$. graph ${\mathbb{G}}=(V,E)$ with positive edge lengths $\{w_{e}\}_{e\in E}$. The continuous graph setting is adopted, i.e. ${\mathbb{G}}$ is regarded as the collection of all nodes in $V$ together with all points forming the edges in $E$. Additionally, ${\mathbb{G}}$ is equipped with the graph metric $d_{{\mathbb{G}}}(x,y)$, which is defined as the length of the shortest path connecting $x$ and $y$ in ${\mathbb{G}}$. We assume that there exists a fixed root node $z_{0}\in V$ such that the shortest path between $z_{0}$ and $x$ is unique for any $x\in{\mathbb{G}}$, i.e., the uniqueness property of the shortest paths (Le et al., 2022).
@@
[…]
@@ 3.1 Sobolev IPM
Sobolev norm. $W^{1,p}({\mathbb{G}},\lambda)$ is a normed space with the Sobolev norm (Adams & Fournier, 2003, §3.1) being defined as
@@
Hereafter, we consider the weight function
$\hat{w}(x):=1+\lambda(\Lambda(x)),\qquad\forall x\in{\mathbb{G}}.$ (5)
@@ 3.2 Regularized Sobolev IPM
We note that in identity (9), both the subgraph $\Lambda(x)$ (Equation (2)) and the weight function $\hat{w}(x)$ (Equation (5)) depend on input point $x$ under the integral over ${\mathbb{G}}$.
For practical applications, we next derive an explicit formula for the integral over graph ${\mathbb{G}}$ in Equation (9) when the input probability measures are supported on nodes $V$ of graph ${\mathbb{G}}$. This gives an efficient method for computing the regularized Sobolev IPM $\hat{{\mathcal{S}}}_{p}$. Note that to achieve this result, we use the length measure on graph ${\mathbb{G}}$ (Le et al., 2022) [^5] See Appendix §B.2 for a review. for the nonnegative Borel measure $\lambda$, i.e., we have $\lambda([x,z])=d_{{\mathbb{G}}}(x,z),\forall x,z\in{\mathbb{G}}$. We summarize the result in the following theorem.
@@
where for each edge $e\in E$ of graph ${\mathbb{G}}$, the number $\beta_{e}$ is given by
$\beta_{e}:=\left\{\begin{array}[]{ll}\log{\left(1+\frac{w_{e}}{1+\lambda(\gamma_{e})}\right)}&\mbox{if }p=2\\ \frac{\left(1+\lambda(\gamma_{e})+w_{e}\right)^{2-p}-\left(1+\lambda(\gamma_{e})\right)^{2-p}}{2-p}&\text{otherwise.}\end{array}\right.$ (11)
@@
Preprocessing. For the computation of the regularized Sobolev IPM $\hat{{\mathcal{S}}}_{p}$ in Equation (10), observe that the set $\gamma_{e}$ (see Equation (2)) and $\beta_{e}$ (see Equation (11)) can be precomputed for all edges $e$ in ${\mathbb{G}}$. Notably, the preprocessing procedure is only involved graph ${\mathbb{G}}$, and it is independent of the input probability measures. As a consequence, we only need to perform this preprocessing procedure once, regardless of the number of input pairs of probability measures that are evaluated by $\hat{{\mathcal{S}}}_{p}$ in applications. More precisely, we apply the Dijkstra algorithm to recompute the shortest paths from root node $z_{0}$ to all other input supports (or vertices) with complexity $\mathcal{O}(|E|+|V|\log{|V|})$ where we write ||$\cdot$|| ⋅ | for the set cardinality. Then, we can evaluate $\gamma_{e}$ and $\beta_{e}$ for each edge $e$ in $E$.
Sparsity of subgraph $\gamma_{e}$ in ${\mathbb{G}}$. Denote $\text{supp}(\mu)$ as the set of supports of measure $\mu$. For any $x\in\text{supp}(\mu)$, its mass is accumulated into $\mu(\gamma_{e})$ if and only if $e\subseteq[z_{0},x]$ (Le et al., 2022). Therefore, let define $E_{\mu,\nu}\subseteq E$ as follows
@@ 4.1 Metric and Its Relations
where ${\mathcal{W}}_{1}$ is the $1$-Wasserstein distance [^8] We review $p$-Wasserstein distance ${\mathcal{W}}_{p}$ in Appendix §B.3. with ground cost $d_{{\mathbb{G}}}$.
@@ 5 Related Works
6 Experiments
In this section, we illustrate the fast computation for the regularized Sobolev IPM, which is comparable to the Sobolev transport (ST), and several-order faster than the standard optimal transport (OT) for measures on a graph. We then show preliminary evidences on the advantages of the regularized Sobolev IPM kernels for document classification and for TDA.
Document classification. We consider $4$ popular document datasets: TWITTER, RECIPE, CLASSIC, AMAZON. The properties of these datasets are summarized in Figure 1. As in Le et al. (2022), we use word2vec word embedding (Mikolov et al., 2013) to map words into vectors in $\mathbb{R}^{300}$. Thus, each document is represented as a probability measure where its supports are in $\mathbb{R}^{300}$, and their corresponding mass is the word frequency in the document.
TDA. We consider orbit recognition on the synthesized Orbit dataset (Adams et al., 2017), and object classification on a $10$-class subset of MPEG7 dataset (Latecki et al., 2000) as in Le et al. (2022). The properties of these datasets are shown in Figure 4. Objects of interest are represented by persistence diagrams (PD), extracted by algebraic topology methods (e.g., persistence homology) (Edelsbrunner & Harer, 2008). Therefore, each PD can be regarded as a distribution of 2D topological feature data points with a uniform mass.
Graph. We consider the graphs ${\mathbb{G}}_{\text{Log}}$ and ${\mathbb{G}}_{\text{Sqrt}}$ (Le et al., 2022, §5) for experiments. Empirically, these graphs satisfy the assumptions in §2, similar to observations in Le et al. (2022). [^10] We review these graphs in Appendix §C.3. We consider $M\!=\!10^{2},10^{3},10^{4}$ as the number of nodes for these graphs, but omit $M=10^{4}$ in MPEG7 due to the small size of this dataset. Recall that graphs ${\mathbb{G}}_{\text{Log}}$ and ${\mathbb{G}}_{\text{Sqrt}}$ have $(M\log{M})$ and $M^{3/2}$ edges respectively.
Root node $z_{0}$. The regularized Sobolev IPM $\hat{{\mathcal{S}}}$ is defined on graph ${\mathbb{G}}$ with root node $z_{0}$. Notice that $z_{0}$ characterizes the graph derivative for functions on ${\mathbb{G}}$ (Definition 3.1). Much as Sobolev transport (Le et al., 2022), we employ the sliced approach to weaken this dependency. More precisely, we uniformly average $\hat{{\mathcal{S}}}$ over different choices of root nodes, i.e., a sliced variant for $\hat{{\mathcal{S}}}$.
@@ 6 Experiments
Classification. We use kernelized support vector machine (SVM) for document classification and TDA. Specifically, we evaluate regularized Sobolev IPM kernels $k_{\hat{\mathcal{S}}_{1}},k_{\hat{\mathcal{S}}_{1.5}^{1.5}},k_{\hat{\mathcal{S}}_{2}^{2}}$; ST kernels $k_{\mathcal{ST}_{1.5}^{1.5}},k_{\mathcal{ST}_{2}^{2}}$, [^11] Note that $k_{\mathcal{ST}_{\!1}}=k_{\hat{\mathcal{S}}_{1}}$ by Proposition 4.3.; and kernels $\exp(-t\bar{d}(\cdot,\cdot))$ with $t>0$, where $\bar{d}$ is a distance (e.g., OT with ground cost $d_{{\mathbb{G}}}$, tree-Wasserstein (TW) with tree sampled from graph ${\mathbb{G}}$) as considered in Le et al. (2022). The kernels corresponding to the two choices of $\bar{d}$ are respectively denoted as $k_{OT},k_{TW}$. Note that kernel $k_{OT}$ is empirically indefinite. We regularize its Gram matrices by adding a sufficiently large diagonal term as in Cuturi (2013).
SVM results and discussions. The reported results for all datasets are with $M=10^{4}$, except for MPEG7 with $M=10^{3}$ due to the small size of this dataset. We illustrate SVM results and time consumption for kernel matrices on document classification for graphs ${\mathbb{G}}_{\text{Log}}$ and ${\mathbb{G}}_{\text{Sqrt}}$ in Figures 1 and 2 respectively. For TDA, we show the results for ${\mathbb{G}}_{\text{Log}}$ and ${\mathbb{G}}_{\text{Sqrt}}$ in Figures 4 and 5 respectively.
Table 1: The number of pairs of measures on datasets for SVM.
@@
Additionally, we illustrate performance of sliced variants with graph ${\mathbb{G}}_{\text{Log}}$ in Figures 3 and 6 for document classification and TDA respectively. When the number of slices increases, their performances are improved but it comes with a trade-off on increasing the computational time linearly.

### v2

@@ Abstract
We investigate the Sobolev IPM problem for probability measures supported on a graph metric space. Sobolev IPM is an important instance of integral probability metrics (IPM), and is obtained by constraining a critic function within a unit ball defined by the Sobolev norm. In particular, it has been used to compare probability measures and is crucial for several theoretical works in machine learning. However, to our knowledge, there are no efficient algorithmic approaches to compute Sobolev IPM effectively, which hinders its practical applications. In this work, we establish a relation between Sobolev norm and weighted $L^{p}$-norm, and leverage it to propose a novel regularization for Sobolev IPM. By exploiting the graph structure, we demonstrate that the regularized Sobolev IPM provides a closed-form expression for fast computation. This advancement addresses long-standing computational challenges, and paves the way to apply Sobolev IPM for practical applications, even in large-scale settings. Additionally, the regularized Sobolev IPM is negative definite. Utilizing this property, we design positive-definite kernels upon the regularized Sobolev IPM, and provide preliminary evidences of their advantages for comparing probability measures on a given graph for document classification and topological data analysis.
Sobolev IPM, probability measures on a graph, scalability, integral probability metrics (IPM), Sobolev norm, weighted Lp-norm
@@ 1 Introduction
Probability measures are widely used to represent objects of interest across various research fields. For instance, in natural language processing, documents can be viewed as distributions over words (Sparck Jones, 1972; Kusner et al., 2015; Le et al., 2019), or distributions over topics (Blei et al., 2003; Yurochkin et al., 2019). In computer vision and graphics, 3D objects are often represented as point clouds, which are distributions of 3D data points (Achlioptas et al., 2018; Hua et al., 2018; Wang et al., 2019; Wu et al., 2019; Nguyen et al., 2021). In topological data analysis (TDA), persistence diagrams (PD) are used to represent complex structural objects (Le & Yamada, 2018; Rieck et al., 2019; Zhao & Wang, 2019; Divol & Lacombe, 2021). Specifically, PDs can be regarded as distributions of 2D data points, where each point characterizes the birth and death time of a particular topological feature (Edelsbrunner & Harer, 2008).
@@
Our contributions are three-fold as follows:
@@
  • We demonstrate that the regularized Sobolev IPM is negative definite. We then design positive definite kernels based on the regularized Sobolev IPM and provide preliminary evidence of their advantages to compare probability measures on a given graph in document classification and topological data analysis (TDA).
@@
In §2, we introduce the notations used throughout our proposals. We detail our novel regularization for Sobolev IPM in §3. In §4, we prove the metric property for the regularized Sobolev IPM and establish its connection to the original Sobolev IPM, ST, and OT on a graph. Additionally, we demonstrate the negative definiteness for the regularized Sobolev IPM and propose positive definite kernels based on it. In §5, we discuss related works. In §6, we empirically illustrate the computational advantages of the regularized Sobolev IPM and provide preliminary evidence of its benefits in document classification and TDA. Finally, in §7, we offer concluding remarks. Proofs for theoretical results and additional materials are deferred to the Appendices. Additionally, we have released code for our proposed approach. [^1] The code repository is on https://github.com/lttam/Sobolev-IPM.
@@ 2 Preliminaries
Figure 1: A geodetic graph with $10$ nodes $\left\{x_{1},x_{2},\dotsc,x_{10}\right\}$ and $15$ edges $\left\{e_{1},e_{2},\dotsc,e_{15}\right\}$, and each edge length equals to $1$, i.e., $w_{e_{j}}=1,\forall j$. For any $x_{i},x_{j}$, there is a unique shortest path between them, with a length $2$. Therefore, it satisfies the uniqueness property of the shortest paths. Let $x_{1}$ be the unique-path root node (i.e., $z_{0}=x_{1}$), and subgraph $\widetilde{{\mathbb{G}}}$ containing $3$ nodes $\left\{x_{4},x_{5},x_{10}\right\}$ and $2$ edges $\left\{e_{4},e_{10}\right\}$, then we have $\Lambda(x_{5})=\gamma(e_{5})=\widetilde{{\mathbb{G}}}$.
Graph.
We follow the same graph setting in Le et al. (2022) for measures and functions. Precisely, let $V,E$ denote the sets of nodes and edges respectively. We consider a connected, undirected, and physical [^2] In the sense that $V$ is a subset of Euclidean space $\mathbb{R}^{n}$; each edge $e\in E$ is the standard line segment in $\mathbb{R}^{n}$ connecting the two vertices of the edge $e$. graph ${\mathbb{G}}=(V,E)$ with positive edge lengths $\{w_{e}\}_{e\in E}$. The continuous graph setting is adopted, i.e. ${\mathbb{G}}$ is regarded as the collection of all nodes in $V$ together with all points forming the edges in $E$. Additionally, ${\mathbb{G}}$ is equipped with the graph metric $d_{{\mathbb{G}}}(x,y)$, which is defined as the length of the shortest path connecting $x$ and $y$ in ${\mathbb{G}}$. We assume that there exists a fixed root node $z_{0}\in V$ such that the shortest path between $z_{0}$ and $x$ is unique for any $x\in{\mathbb{G}}$, i.e., the uniqueness property of the shortest paths (Le et al., 2022).  [^3] See Figure 1 for an illustration of a geodetic graph which satisfies the uniqueness property of the shortest paths.
@@
See Figure 1 for an illustration of these graph notions.
Measures and functions on graph.
Let ${\mathcal{P}}({\mathbb{G}})$ (resp. ${\mathcal{P}}({\mathbb{G}}\times{\mathbb{G}})$) denote the set of all nonnegative Borel measures on ${\mathbb{G}}$ (resp. ${\mathbb{G}}\times{\mathbb{G}}$) with a finite mass.
@@ 3.1 Sobolev IPM
Sobolev norm.
$W^{1,p}({\mathbb{G}},\lambda)$ is a normed space with the Sobolev norm (Adams & Fournier, 2003, §3.1) being defined as
@@
To our knowledge, there are no efficient algorithmic approaches to compute the Sobolev IPM in Equation (4) effectively, especially for large-scale settings.
@@
Hereafter, we consider the weight function, defined as
$\hat{w}(x):=1+\lambda(\Lambda(x)),\qquad\forall x\in{\mathbb{G}},$ (5)
where recall that $\Lambda(x)$ is a subgraph (see Equation (2), and Figure 1), and $\lambda(\cdot)$ is measure $\lambda$ on a set. For instance, when $\lambda$ is the length measure on graph ${\mathbb{G}}$ (Le et al., 2022) [^6] See Appendix §B.2 for a review., then $\lambda(\Lambda(x))$ is the total length of $\Lambda(x)$.
@@ 3.2 Regularized Sobolev IPM
Notice that in identity (9), both the subgraph $\Lambda(x)$ (Equation (2)) and the weight function $\hat{w}(x)$ (Equation (5)) depend on input point $x$ under the integral over ${\mathbb{G}}$.
For practical applications, we next derive an explicit formula for the integral over graph ${\mathbb{G}}$ in Equation (9) when the input probability measures are supported on nodes $V$ of graph ${\mathbb{G}}$. This gives an efficient method for computing the regularized Sobolev IPM $\hat{{\mathcal{S}}}_{p}$. Note that to achieve this result, we use the length measure on graph ${\mathbb{G}}$ (Le et al., 2022) for the nonnegative Borel measure $\lambda$, i.e., we have $\lambda([x,z])=d_{{\mathbb{G}}}(x,z),\forall x,z\in{\mathbb{G}}$. We summarize the result in the following theorem.
@@
where for each edge $e\in E$ of graph ${\mathbb{G}}$, the scalar number $\beta_{e}$ is given by
$\beta_{e}:=\left\{\begin{array}[]{ll}\log{\left(1+\frac{w_{e}}{1+\lambda(\gamma_{e})}\right)}&\mbox{if }p=2,\\ \frac{\left(1+\lambda(\gamma_{e})+w_{e}\right)^{2-p}-\left(1+\lambda(\gamma_{e})\right)^{2-p}}{2-p}&\text{otherwise.}\end{array}\right.$ (11)
@@
Preprocessing.
For the computation of the regularized Sobolev IPM $\hat{{\mathcal{S}}}_{p}$ in Equation (10), observe that the set $\gamma_{e}$ (see Equation (2)) and $\beta_{e}$ (see Equation (11)) can be precomputed for all edges $e$ in ${\mathbb{G}}$. Notably, the preprocessing procedure is only involved graph ${\mathbb{G}}$, and it is independent of the input probability measures. Consequently, we only need to perform this preprocessing procedure once, regardless of the number of input pairs of probability measures that are evaluated by the regularized Sobolev IPM $\hat{{\mathcal{S}}}_{p}$ in applications. More precisely, we apply the Dijkstra algorithm to recompute the shortest paths from root node $z_{0}$ to all other input supports (or vertices) with complexity $\mathcal{O}(|E|+|V|\log{|V|})$ where we write ||$\cdot$|| ⋅ | for the set cardinality. Then, we can evaluate $\gamma_{e}$ and $\beta_{e}$ for each edge $e$ in $E$.
Sparsity of subgraph $\gamma_{e}$ in ${\mathbb{G}}$.
Denote $\text{supp}(\mu)$ as the set of supports of measure $\mu$. For any $x\in\text{supp}(\mu)$, its mass is accumulated into $\mu(\gamma_{e})$ if and only if $e\subseteq[z_{0},x]$ (Le et al., 2022). Therefore, let define $E_{\mu,\nu}\subseteq E$ as follows
@@
Following the preprocessing procedure and the sparsity of subgraph $\gamma_{e}$ for any edge $e$ in ${\mathbb{G}}$, we only need to check the precomputed shortest paths from root node $z_{0}$ to each support of input measures $\mu,\nu$ to respectively obtain $\mu(\gamma_{e}),\nu(\gamma_{e})$ for all edges $e\in E_{\mu,\nu}$ in graph ${\mathbb{G}}$. Additionally, $\beta_{e}$ is also precomputed for all edge $e$ in ${\mathbb{G}}$. Therefore, it is straightforward to implement the computation of $\hat{{\mathcal{S}}}_{p}(\mu,\nu)$ in Equation (10) with complexity $\mathcal{O}(|E_{\mu,\nu}|)$.
@@ 4.1 Metric and Its Relations
where ${\mathcal{W}}_{1}$ is the $1$-Wasserstein distance [^10] We review $p$-Wasserstein distance ${\mathcal{W}}_{p}$ in Appendix §B.3. with ground cost $d_{{\mathbb{G}}}$ (i.e., tree-Wasserstein distance (Le et al., 2019)).
@@ 5 Related Works
6 Experiments
In this section, we illustrate the fast computation for the regularized Sobolev IPM, which is comparable to the Sobolev transport (ST), and several-order faster than the standard optimal transport (OT) for measures on a graph. We then show preliminary evidences on the advantages of the regularized Sobolev IPM kernels to compare probability measures on a given graph under the same settings for document classification and for TDA.
Document classification.
We consider $4$ popular document datasets: TWITTER, RECIPE, CLASSIC, AMAZON. The properties of these datasets are summarized in Figure 2. As in Le et al. (2022), we use word2vec word embedding (Mikolov et al., 2013) to map words into vectors in $\mathbb{R}^{300}$. Thus, each document is represented as a probability measure where its supports are in $\mathbb{R}^{300}$, and their corresponding mass is the word frequency in the document.
TDA.
We consider orbit recognition on the synthesized Orbit dataset (Adams et al., 2017), and object classification on a $10$-class subset of MPEG7 dataset (Latecki et al., 2000) as in Le et al. (2022). The properties of these datasets are shown in Figure 5. Objects of interest are represented by persistence diagrams (PD), extracted by algebraic topology methods (e.g., persistence homology) (Edelsbrunner & Harer, 2008). Therefore, each PD can be regarded as a distribution of 2D topological feature data points with a uniform mass.
Graph.
We consider the graphs ${\mathbb{G}}_{\text{Log}}$ and ${\mathbb{G}}_{\text{Sqrt}}$ (Le et al., 2022, §5) for experiments. Empirically, these graphs satisfy the assumptions in §2, similar to observations in Le et al. (2022). [^12] We review these graphs in Appendix §C.3. We consider $M\!=\!10^{2},10^{3},10^{4}$ as the number of nodes for these graphs, but omit $M=10^{4}$ in MPEG7 due to the small size of this dataset. Recall that graphs ${\mathbb{G}}_{\text{Log}}$ and ${\mathbb{G}}_{\text{Sqrt}}$ have $(M\log{M})$ and $M^{3/2}$ edges respectively.
Table 1: The number of pairs of probability measures on datasets for kernelized SVM.
@@ 6 Experiments
Root node $z_{0}$.
The regularized Sobolev IPM $\hat{{\mathcal{S}}}$ is defined on graph ${\mathbb{G}}$ with root node $z_{0}$. Notice that $z_{0}$ characterizes the graph derivative for functions on ${\mathbb{G}}$ (Definition 3.1). Much as Sobolev transport (Le et al., 2022), we employ the sliced approach to weaken this dependency. More precisely, we uniformly average $\hat{{\mathcal{S}}}$ over different choices of root nodes, i.e., a sliced variant for $\hat{{\mathcal{S}}}$.
Classification.
We use kernelized support vector machine (SVM) for document classification and TDA. Specifically, we evaluate regularized Sobolev IPM kernels $k_{\hat{\mathcal{S}}_{1}},k_{\hat{\mathcal{S}}_{1.5}^{1.5}},k_{\hat{\mathcal{S}}_{2}^{2}}$; ST kernels $k_{\mathcal{ST}_{1.5}^{1.5}},k_{\mathcal{ST}_{2}^{2}}$, [^13] Note that $k_{\mathcal{ST}_{\!1}}=k_{\hat{\mathcal{S}}_{1}}$ by Proposition 4.3.; and kernels $\exp(-t\bar{d}(\cdot,\cdot))$ with $t>0$, where $\bar{d}$ is a distance (e.g., OT with ground cost $d_{{\mathbb{G}}}$, tree-Wasserstein (TW) with tree sampled from graph ${\mathbb{G}}$) as considered in Le et al. (2022). The kernels corresponding to the two choices of $\bar{d}$ are respectively denoted as $k_{OT},k_{TW}$. Note that kernel $k_{OT}$ is empirically indefinite. We regularize its Gram matrices by adding a sufficiently large diagonal term as in Cuturi (2013).
SVM results and discussions.
The reported results for all datasets are with $M=10^{4}$, except for MPEG7 with $M=10^{3}$ due to the small size of this dataset. We illustrate SVM results and time consumption for kernel matrices on document classification for graphs ${\mathbb{G}}_{\text{Log}}$ and ${\mathbb{G}}_{\text{Sqrt}}$ in Figures 2 and 3 respectively. For TDA, we show the results for ${\mathbb{G}}_{\text{Log}}$ and ${\mathbb{G}}_{\text{Sqrt}}$ in Figures 5 and 6 respectively.
@@
Additionally, we illustrate performance of sliced variants with graph ${\mathbb{G}}_{\text{Log}}$ in Figures 4 and 7 for document classification and TDA respectively. When the number of slices increases, their performances are improved but it comes with a trade-off on increasing the computational time linearly.